A wedge shaped film has refractive index 1.34 and thickness of extreme sides 0 (zero) and t. If a light of wavelength 492 nm is incident normally on it and 20 fringes are obtained, determine t.
1
Expert's answer
2020-02-17T09:10:47-0500

As per the given data in the question,

Refractive index of the film=1.34

thickness of the extreme side 0 and t

wavelength of the light (λ)=492nm(\lambda)=492nm

(λ)=492nm

Number of fringes=20

As per the rule of x-ray reflection on the single layer

"2t\\sqrt{n^2\u2212\\sin^2i}-\\dfrac{\\lambda}{2}\u200b=m\u03bb"

n=refractive index of the medium

Light is getting incident normally, so i=0i=0

i=0

So,

"2tn-\dfrac{\lambda}{2}=m\lambda"

"\Rightarrow t=\dfrac{m\lambda+\dfrac{\lambda}{2}}{2n}"

"\Rightarrow t=\dfrac{20\times492\times10^{-9}+246\times 10^{-9}}{2\times 1.34}"


"\Rightarrow t=9931.8\times10^{-9}m"


"\Rightarrow t=9.9318\times 10^{-6}m"

"t=9.93\mu m"